ABO Blood Type allele frequencies:

[I^O] = 0.7 , [I^A] = 0.2 , [I^B] = 0.1

For the Rh factor frequencies:

[Rh+] = 0.8 , [Rh-] = 0.2

1. In this population what is the frequency of the type "O", Rh positive individuals?

2. The rarest type of blood is type AB Rh negative. How many of these are expected in 1000 individuals?

3. In this population what will be the frequency go type "A", Rh negative individuals?

4. Assume the allele frequencies for Rh alleles above. What will be the allele frequency of Rh- after one generation of selecting this population against the Rh- allele. Assume W22 = 0.
Answer #1

1. Here the frequency of O allele is 0.7, therefore frequency of indivuals having o group= (0.7)2= 0.49

Frequency of Rh positive allele is 0.8, so frequency of individuals with Rh positive will be p2+2pq= 0.64+0.32= 0.96

Now, frequency of o group with Rh positive individuals= 0.49X 0.96= 0.47.04. Answer.

2. Frequency of A allele is 0.2 and frequency of B allele is 0.1, so frequency of individuals with AB group will be 2pq= 2X0.2X0.1= 0.04.

Frequency of Rh negative allele is 0.2, so frequency of individuals with Rh negative= 0.2X0.2= 0.04.

Therefore, frequency of AB with Rh negative individuals= 0.04X0.04= 0.0016

So in 1000 individuals, 0.0016X 1000= 1.6 individuals will be AB with Rh negative.

3. Frequency of A allele is 0.2 and frequency of O allele is 0.7, so frequency of individuals with A group will be= 0.04+0.28= 0.32.

frequency of individuals with Rh negative= 0.04

Therefore, frequency of individuals with A and Rh will be 0.32X0.04= 0.0128 Answer.